# L2M-Reg: Building-level Uncertainty-aware Registration of Outdoor LiDAR Point Clouds and Semantic 3D City Models

### v2

@@ 4.4.2 Quantitative Comparison
In this section, the aforementioned evaluation metrics are used to quantitatively compare the four valid registration methods (TriICP, GICP, PLADE, and L2M-Reg) based on the check points manually selected from stable areas. Table 2 shows the number of check points used in each dataset. In the Street Building dataset, all ground points in the central areas of the road were used for vertical error evaluation, considering local layering issues in the road model.
Table 2: The Number of Check Points Used in Three Datasets
TUM0501 Building Pinakothek Street Building
Horizontal check points 12 22 16
Vertical check points 8 20 620
@@
Table 3 reports the average registration error in both horizontal and vertical directions. L2M-Reg achieves the best performance on all three datasets, which is consistent with the qualitative results discussed before. The sole exception lies in the vertical error of the Street Building dataset, where L2M-Reg (2.30 cm) performs slightly worse than GICP (1.89 cm). This minor discrepancy is primarily attributed to the reference ground data, which is derived from an additional road model rather than a standard DTM. Minor layering artifacts in this road model introduce local inconsistencies that bias the vertical error calculation. Nevertheless, the resulting loss (less than 0.5 cm) is negligible for most practical purposes.
Table 3: Average Error Comparison on Three Datasets (cm)
TriICP GICP PLADE L2M-Reg (Ours)
TUM0501 $\mathit{Err}_{\mathrm{H}}$ 14.07 5.75 0.98
$\mathit{Err}_{\mathrm{V}}$ 24.44 6.24 1.20
Pinakothek $\mathit{Err}_{\mathrm{H}}$ 12.23 18.13 32.36 3.01
$\mathit{Err}_{\mathrm{V}}$ 24.86 9.20 31.13 2.03
Street Building $\mathit{Err}_{\mathrm{H}}$ 24.31 6.68 11.77 4.87
$\mathit{Err}_{\mathrm{V}}$ 2.89 1.89 6.71 2.30
@@
Note: “–” indicates that the method failed to produce valid results; Bold values indicate the best performance.
@@
Table 4 presents the standard deviation of the M3C2 distance across all check points. L2M-Reg consistently achieves the lowest standard deviation across all three datasets, indicating its effectiveness in producing globally optimal and stable registration results. A similar issue can be observed again in the vertical direction of the Street Building dataset. As previously noted, this slight inconsistency primarily stems from the characteristics of the road model. Given the magnitude of the difference, the deviation remains within a negligible range.
Table 4: Standard Deviation Comparison on Three Datasets (cm)
TriICP GICP PLADE L2M-Reg (Ours)
TUM0501 $\mathit{Std}_{\mathrm{H}}$ 7.46 3.39 0.81
$\mathit{Std}_{\mathrm{V}}$ 3.24 3.71 0.72
Pinakothek $\mathit{Std}_{\mathrm{H}}$ 5.15 6.83 17.09 3.91
$\mathit{Std}_{\mathrm{V}}$ 2.22 1.32 4.16 0.72
Street Building $\mathit{Std}_{\mathrm{H}}$ 22.3 6.06 9.99 6.05
$\mathit{Std}_{\mathrm{V}}$ 2.19 1.37 3.28 1.50
@@
Note: “–” indicates that the method failed to produce valid results; Bold values indicate the best performance.
@@
In summary, both qualitative and quantitative comparisons demonstrate that L2M-Reg achieves superior performance on both the TUM0501 and Pinakothek datasets, as reflected in lower average error and standard deviations. For the Street Building dataset, L2M-Reg also outperforms other methods in horizontal error. Although it does not exhibit a clear advantage in vertical error for this dataset, its performance remains fully comparable to that of the best-performing method (GICP). Overall, the results prove the effectiveness of L2M-Reg and highlight its leading performance across diverse building scenarios.
@@ 4.4.3 Efficiency Analysis
Table 5 summarizes the running times of L2M-Reg in comparison to GICP, TriICP and PLADE on three datasets. All methods are evaluated under identical hardware conditions, with a maximum iteration limit of 100. Scantra is excluded from this comparison as it failed to yield valid results.
The results in Table 5 demonstrate that L2M-Reg exhibits promising computational efficiency, particularly on the Pinakothek and Street Building datasets, which contain larger numbers of points. The primary reason is that both ICP-based method and PLADE operate on the entire point clouds, leading to computational times scaling linearly with the number of points. In contrast, L2M-Reg significantly reduces the computing time by first fixing correspondence relationships within data association and selectively processing points from building plinth. Moreover, as a plane-based method, its complexity is less sensitive to point density, leading to notable improvements in runtime performance.
Table 5: Comparison of Computational Efficiency (Bold values indicate the best performance)
TUM0501 Building Pinakothek Street Building
GICP 30.4s 329.5s 328.6s
TriICP 38.5s 356.1s 120.2s
PLADE 81.2s 182.4s 140.9s
L2M-Reg (Ours) 38.0s 135.9s 52.7s
@@ 5 Discussion
This part further discusses the advantages of L2M-Reg in Section 5.1, comparison among three plane-based methods in Section 5.2, reasons of using road model in the Street Building Dataset in Section 5.3, and limitations of L2M-Reg in Section 5.4, providing a deeper understanding of its performance and applicability.
@@ 5.1 Advantages of Using Existing Model Semantics and Pseudo-plane Constraint
Compared to existing ICP-based (e.g., GICP and TriICP) and plane-based (e.g., PLADE and Scantra) registration methods that exclusively process point clouds generated from LoD2 models, L2M-Reg uniquely leverages built-in semantic information in LoD2 models. Using these semantic attributes and distinct plane identifiers, facade elements can be efficiently and accurately extracted, enabling precise plane correspondence establishment. This strategy bypasses the conventional two‑step workflow of feature extraction followed by correspondence matching. Instead, it directly streamlines correspondence identification by leveraging the existing semantic information in LoD2 models, thus improving the registration robustness and efficiency.
As one of the key innovations of L2M-Reg, introducing the pseudo-plane constraint offers three distinct advantages. First, it enables 2D–3D decoupled parameter estimation, effectively preventing low-quality elevation data from influencing the estimation of other parameters. Second, it is formulated within a clear and interpretable mathematical framework under the GHM and can be implemented with minimal complexity. Third, compared to the alternative of manually fixing $t_{z}$ to zero in GHM to isolate vertical errors, the pseudo-plane strategy provides greater flexibility and adaptability, allowing for extension to other facade orientations. Moreover, in scenarios involving sloped or uneven ground surfaces, directly constraining $t_{z}$ in GHM may introduce additional errors, further highlighting the advantages of the pseudo-plane strategy.
@@ 5.2 Interpretation of Performance Differences among Three Plane-based Methods
The design of PLADE and Scantra lies in extracting a large number of planes prior to establishing correspondences using geometric descriptors. While effective in general urban scenes, this strategy proves unsuitable for individual buildings, which typically contain a limited number of planes with diverse orientations. In addition, their correspondence establishment heavily rely on rich geometric structures and distinctive features. This makes them less effective when applied to geometrically simplified or feature-sparse LoD2 models of individual buildings, accounting for their performance in the evaluated scenarios.
5.3 Explanation of Using Road Model in the Street Building Dataset
@@
In the Street Building dataset, a local road model is used instead of DTM to provide constraints in the vertical direction. This choice is motivated by two main factors. First, increasingly detailed ground and road models—such as high-definition maps for autonomous driving—are becoming available for local areas. Although their spatial coverage is currently limited, these datasets often offer higher accuracy than DTM derived from ALS and can serve as more reliable sources of absolute elevation. Second, the effectiveness of L2M-Reg when using DTM data has already been demonstrated based on the TUM0501 and Pinakothek datasets. Owing to the 2D–3D decoupled parameter estimation strategy introduced in this study, the estimation of vertical translation parameters does not affect other transformation parameters. Therefore, using road model in the Street Building dataset further illustrates the adaptability of L2M-Reg.
@@
While a slight difference in vertical accuracy is observed between L2M-Reg and the best-performing baseline (GICP) in the quantitative results (see Tables 3 and 4), this discrepancy is primarily attributed to redundant layers within the used road model, which distort the accuracy evaluation. Provided higher-quality road data in the future, L2M-Reg can readily adapt to such data sources.

### v3

@@ 4.4.2 Quantitative Comparison
In this section, the aforementioned evaluation metrics are used to quantitatively compare the five valid registration methods (KISS-ICP, TriICP, GICP, PLADE, and L2M-Reg) based on the check points manually selected from stable areas. Table 2 shows the number of check points used in each dataset. In the Street Building dataset, all ground points in the central areas of the road were used for vertical error evaluation, considering local layering issues in the road model.
Table 2: The Number of Check Points Used in Five Datasets
Horizontal check points Vertical check points
TUM0501 Building 12 8
Pinakothek 22 20
Street Building 16 620
Restaurant 12 12
Ingolstadt Store 10 10
@@
Table 3 reports the average registration error in both horizontal and vertical directions. L2M-Reg achieves the best performance on all datasets, which is consistent with the qualitative results discussed before. There are only two exceptions regarding the vertical error of the Street Building and Restaurant datasets, where L2M-Reg performs slightly worse. This minor discrepancy is primarily attributed to the reference ground data, which is derived from an additional road model rather than a standard DTM for the Street Building. Minor layering artifacts in this road model introduce local inconsistencies that bias the vertical error calculation. Nevertheless, the resulting loss (around 0.5 cm) is negligible for most practical purposes.
Table 3: Average Error Comparison on Five Datasets (cm)
KISS-ICP TriICP GICP PLADE L2M-Reg (Ours)
TUM0501 Building $\mathit{Err}_{\mathrm{H}}$ 8.68 14.07 5.75 0.98
$\mathit{Err}_{\mathrm{V}}$ 6.74 24.44 6.24 1.20
Pinakothek $\mathit{Err}_{\mathrm{H}}$ 26.75 12.23 18.13 32.36 3.01
$\mathit{Err}_{\mathrm{V}}$ 2.82 24.86 9.20 31.13 2.03
Street Building $\mathit{Err}_{\mathrm{H}}$ 9.18 24.31 6.68 11.77 4.87
$\mathit{Err}_{\mathrm{V}}$ 2.80 2.89 1.89 6.71 2.30
Restaurant $\mathit{Err}_{\mathrm{H}}$ 9.22 5.75 5.90 5.29
$\mathit{Err}_{\mathrm{V}}$ 1.36 1.92 1.64 1.91
Ingolstadt Store $\mathit{Err}_{\mathrm{H}}$ 11.97 13.16 4.44 1.23
$\mathit{Err}_{\mathrm{V}}$ 4.16 1.17 1.24 0.78
@@
Note: ”–” indicates that the method failed to produce valid results; Bold values indicate the best results.
@@
Table 4 presents the standard deviation of the M3C2 distance across all check points. L2M-Reg consistently achieves the lowest standard deviation across all datasets except the Restaurant, indicating its effectiveness in producing globally optimal and stable registration results. A similar issue can be observed again in the vertical direction of the Street Building dataset. As previously noted, this slight inconsistency primarily stems from the characteristics of the road model. While L2M-Reg performs slightly worse than the best-performing method on the Restaurant dataset, the differences are all below 0.5 cm and are therefore considered acceptable.
Table 4: Standard Deviation Comparison on Five Datasets (cm)
KISS-ICP TriICP GICP PLADE L2M-Reg (Ours)
TUM0501 Building $\mathit{Std}_{\mathrm{H}}$ 3.44 7.46 3.39 0.81
$\mathit{Std}_{\mathrm{V}}$ 5.08 3.24 3.71 0.72
Pinakothek $\mathit{Std}_{\mathrm{H}}$ 12.17 5.15 6.83 17.09 3.91
$\mathit{Std}_{\mathrm{V}}$ 1.65 2.22 1.32 4.16 0.72
Street Building $\mathit{Std}_{\mathrm{H}}$ 8.51 22.3 6.06 9.99 6.05
$\mathit{Std}_{\mathrm{V}}$ 2.05 2.19 1.37 3.28 1.50
Restaurant $\mathit{Std}_{\mathrm{H}}$ 6.92 4.63 4.62 4.94
$\mathit{Std}_{\mathrm{V}}$ 0.96 1.06 0.94 1.33
Ingolstadt Store $\mathit{Std}_{\mathrm{H}}$ 3.24 4.17 2.57 1.21
$\mathit{Std}_{\mathrm{V}}$ 1.51 0.76 0.78 0.64
@@
Note: “–” indicates that the method failed to produce valid results; Bold values indicate the best results.
@@
In summary, both qualitative and quantitative comparisons demonstrate that L2M-Reg achieves superior performance, as reflected in lower average error and standard deviations. For the Street Building and Restaurant datasets, although it does not exhibit a clear advantage, its performance remains fully comparable to that of the best-performing method. Overall, these comparisons prove the effectiveness of L2M-Reg and highlight its leading performance across diverse building scenarios.
@@ 4.4.3 Efficiency Analysis
Table 5 summarizes the running times of L2M-Reg in comparison to KISS-ICP, GICP, TriICP, and PLADE on five datasets. All methods are evaluated under identical hardware conditions, with a maximum iteration limit of 100. Scantra is excluded from this comparison as it failed to yield valid results.
The results in Table 5 demonstrate that L2M-Reg exhibits promising computational efficiency, particularly on the Pinakothek and Restaurant datasets, which contain larger numbers of points. The primary reason is that GICP, TriICP, and PLADE operate on the entire point clouds, leading to computational times scaling linearly with the number of points. In contrast, L2M-Reg significantly reduces the computing time by first fixing correspondence relationships within data association and selectively processing points from the building plinth. Moreover, as a plane-based method, its complexity is less sensitive to point density, leading to notable improvements in runtime performance.
Additionally, compared to KISS-ICP, L2M-Reg achieves superior accuracy as shown in Table 3, but loses its lead in computational efficiency. This is because KISS-ICP employs very aggressive downsampling in its design to accelerate processing. Such a design primarily serves real-time demanding scenarios like autonomous driving and robotics, thereby sacrificing certain accuracy for improved efficiency.
Table 5: Comparison of Computational Efficiency
GICP TriICP KISS-ICP PLADE L2M-Reg (Ours)
TUM0501 Building 30.4s 38.5s 5.7s 81.2s 38.0s
Pinakothek 329.5s 356.1s 18.5s 182.4s 135.9s
Street Building 328.6s 120.2s 10.8s 140.9s 52.7s
Restaurant 1221.0s 328.5s 31.8s 167.5s 88.6s
Ingolstadt Store 258.0s 24.5s 11.2s 80.5s 22.0s
Note: Best results are highlighted in bold, and the second best results are underlined.
@@ 5 Discussion
This part further discusses the advantages of L2M-Reg in Section 5.1, comparison among three plane-based methods in Section 5.2, and reasons for using the road model in some datasets in Section 5.3. Moreover, Section 5.4, 5.5, and 5.6 give further information regarding the ablation study of GC check, robustness to different initial poses, and generalizability and prospective practical applications of L2M-Reg, respectively, providing a deeper understanding of its performance and applicability. Current limitations of this study are discussed in Section 5.7.
@@ 5.1 Advantages of Using Existing Model Semantics and Pseudo-plane Constraint
Compared to existing ICP-based (e.g., KISS-ICP, GICP, and TriICP) and plane-based (e.g., PLADE and Scantra) registration methods that exclusively process point clouds generated from LoD2 models, L2M-Reg uniquely leverages built-in semantic information in LoD2 models. Using these semantic attributes and distinct plane identifiers, facade elements can be efficiently and accurately extracted, enabling precise plane correspondence establishment. This strategy bypasses the conventional two‑step workflow of feature extraction followed by correspondence matching. Instead, it directly streamlines correspondence identification by leveraging the existing semantic information in LoD2 models, thus improving the registration robustness and efficiency.
As one of the key innovations of L2M-Reg, introducing the pseudo-plane constraint offers three distinct advantages. First, it enables 2D-3D decoupled parameter estimation, effectively preventing low-quality elevation data from influencing the estimation of other parameters. Second, it is formulated within a clear and interpretable mathematical framework under the GHM and can be implemented with minimal complexity. Third, compared to the alternative of manually fixing $t_{z}$ to zero in GHM to isolate vertical errors, the pseudo-plane strategy provides greater flexibility and adaptability, allowing for extension to other facade orientations. Moreover, in scenarios involving sloped or uneven ground surfaces, directly constraining $t_{z}$ in GHM may introduce additional errors, further highlighting the advantages of the pseudo-plane strategy.
@@ 5.2 Interpretation of Performance Differences among Three Plane-based Methods
The design of PLADE and Scantra lies in extracting a large number of planes prior to establishing correspondences using geometric descriptors. While effective in general urban scenes, this strategy proves unsuitable for individual buildings, which typically contain a limited number of planes with diverse orientations. In addition, their correspondence establishment heavily relies on rich geometric structures and distinctive features. This makes them less effective when applied to geometrically simplified or feature-sparse LoD2 models of individual buildings, accounting for their performance in the evaluated scenarios.
5.3 Explanation of Using Road Model
@@
In the Street Building and Ingolstadt Store datasets, two local road models produced by different agencies are used instead of the DTM to provide constraints in the vertical direction. This choice is motivated by two main factors. First, increasingly detailed ground and road models, such as high-definition maps for autonomous driving, are becoming available for local areas. Although their spatial coverage is currently limited, these datasets often offer higher accuracy than DTM derived from ALS and can serve as more reliable sources of absolute elevation. Second, the effectiveness of L2M-Reg when using DTM data has already been demonstrated based on the TUM0501 Building, Pinakothek, and Restaurant datasets. Owing to the 2D-3D decoupled estimation strategy introduced in this study, the estimation of vertical translation parameters does not affect other transformation parameters. Therefore, using the road model in the other two datasets further illustrates the adaptability of L2M-Reg.
@@
While a slight difference in vertical accuracy is observed between L2M-Reg and the best-performing baseline (GICP) in the quantitative results (see Tables 3 and 4), this discrepancy is primarily attributed to redundant layers within the used road model in the Street Building dataset, which distort the accuracy evaluation. When higher-quality road data are available, such as in the Ingolstadt Store dataset, L2M-Reg can readily adapt to these data sources and provide accurate estimations. This also demonstrates that L2M-Reg can be applied to data generated by different agencies across various cities.
5.4 Ablation Study on the Geometric Consistency Check
In real-world LiDAR acquisition and processing, point clouds may exhibit layering effects caused by trajectory estimation errors in MLS or multi-station registration errors in TLS. This effect is particularly evident on building facades and can potentially affect plane extraction. When facade point clouds contain multiple layers, conventional RANSAC-based methods tend to extract planes from a single dominant layer and discard the remaining data by default, leading to incomplete or biased plane estimates.
To address this issue, the Geometric Consistency (GC) check is incorporated to improve the robustness of the proposed plane extraction algorithm. It introduces an additional angle-constrained check that enables the reuse of compatible points from adjacent layers. This design preserves the inherent noise robustness of RANSAC while explicitly mitigating performance degradation caused by point cloud layering, thereby enhancing the reliability of the extracted planes.
To further discuss the effectiveness of the GC check, a wall from the TUM0501 Building was selected for controlled analysis. Since L2M-Reg is a plane-based solution, the accuracy of the final result is actually tied to the accuracy of the estimated plane normals. Therefore, different levels of point cloud layering and misalignment were simulated on the original data. Five study cases were designed with progressively increasing inter-layer misalignment angles of $0.5^{\circ}$, $1.0^{\circ}$, $1.5^{\circ}$, $2.0^{\circ}$, and $2.5^{\circ}$.
Figure 14: Absolute Angular Deviation (AAD) with and without GC check, and the corresponding improvement benefiting from GC check in TUM0501 Building. Cases 1 to 5 correspond to inter-layer misalignment angles of $0.5^{\circ}$, $1.0^{\circ}$, $1.5^{\circ}$, $2.0^{\circ}$, and $2.5^{\circ}$, respectively.
The results shown in Figure 14, compare the Absolute Angular Deviation (AAD) of the extracted plane normals from the reference values with and without the GC check, abbreviated as $\mathit{AAD_{GC}}$ and $\mathit{AAD_{noGC}}$, respectively. In addition, the corresponding improvement percentage benefiting from the GC check is defined in Eq. (18):
$\mathit{Improvement_{GC}}=\left(1-\frac{\mathit{AAD_{GC}}}{\mathit{AAD_{noGC}}}\right)\times 100\%.$ (18)
Although the improvement varies across cases, it typically ranges from 20% to 90%, demonstrating the substantial contribution of the GC check to the accuracy of plane extraction under layered point cloud conditions.
5.5 Robustness to Different Initial Poses from Coarse Registration
Additional experiments were conducted to examine the convergence behavior of L2M-Reg under varying levels of coarse registration errors. The TUM0501 Building was used as a representative example, and different rotations and translations were imposed on the original data to simulate increasing degrees of coarse registration errors.
In total, six study cases were designed. Case 1 serves as the baseline, while Cases 2 through 6 apply progressively larger transformations, as summarized in Table 6. The magnitude of the simulated coarse registration errors increases across cases, as quantified by Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE). The registration results obtained by L2M-Reg for all cases are shown in Figure 15. The results indicate that L2M-Reg maintains robust performance and consistently achieves stable fine registration results, even when the RMSE or MAE after coarse registration approaches 1 m. In practice, coarse registration within a 1 m range is readily achievable using existing methods (Diakite and Zlatanova, 2020; Sheik et al., 2022a, b). Therefore, although L2M-Reg requires an initial coarse registration, it remains applicable to most practical scenarios.
Table 6: Different Imposed Transformations on Original Pose
Study Case Imposed Transformations
TUM0501 Building 1 Original
2 Original + $R_{z}(5^{\circ}$)
3 Original + $R_{z}(10^{\circ}$)
4 Original + $R_{z}(15^{\circ}$)
5 Original + $R_{z}(15^{\circ}$) + $t_{z}(0.5m$)
6 Original + $R_{z}(20^{\circ}$) + $t_{z}(1.0m$)
Figure 15: Fine registration results from L2M-Reg under different coarse registration errors. The study case indices from 1 to 6 correspond to the imposed transformations introduced in Table 6.
Furthermore, to examine the convergence boundary of L2M-Reg under varying coarse registration errors more comprehensively, 12 additional test cases were constructed based on Case 6 in Table 6 and Figure 15. All tests used the TUM0501 Building dataset. Larger imposed transformations were applied to simulate more severe coarse registration errors. In addition, multiple initial parameter settings within the GHM were considered. For each level of coarse registration error, four sets of initial parameters were randomly assigned for realistic processing conditions. The results, summarized in Table 7, show that when the coarse registration RMSE is approximately 1 m, L2M-Reg consistently achieves high accuracy regardless of initialization. Similar behavior appears as the RMSE approaches 2 m (1.90 m), with no noticeable loss in convergence accuracy. However, when the initial RMSE increases to around 2.5 m, L2M-Reg’s performance degrades. The average RMSE rises from 0.15 m to 0.21 m.
As an empirical rule of thumb derived from these experiments, L2M-Reg produces reliable results when the initial RMSE is within 2 m. Although convergence may still occur under larger errors, registration accuracy degrades and should be carefully considered in practice. This empirical rule of thumb is also influenced by other factors such as building size, point cloud quality, and structural complexity, and should therefore be adapted to specific scenarios. A more comprehensive investigation of coarse registration and its associated error characteristics is an important direction for future work but falls beyond the scope of this study. Nonetheless, the experiments demonstrate that L2M-Reg consistently delivers reliable performance when the initial RMSE is below 1 m, a level readily achievable by many existing methods.
Table 7: Impact of Different Coarse Registration Errors on L2M-Reg Performance
Dataset ImposedTrans CoarseRegErr before L2M-Reg InitialPara in GHM MeanRegErr after L2M-Reg
$(R_{x},R_{y},R_{z},t_{x},t_{y},t_{z})$ RMSE[m] MAE[m] $(R_{x},R_{y},R_{z},t_{x},t_{y},t_{z})$ RMSE[m] MAE[m]
TUM0501 Building $(0,0,20,0,0,1)$ 0.99 0.82 $(0,0,0,0,0,0)$ 0.14 0.12
$(0,0,5,0,0,0.5)$
$(0,0,10,0,0,0.8)$
$(10,10,10,0.5,0.5,0.5)$
$(0,0,20,0,0,2)$ 1.90 1.72 $(0,0,0,0,0,0)$ 0.15 0.13
$(0,0,5,0,0,0.5)$
$(0,0,10,0.5,0.5,1.5)$
$(10,10,0,1,1,1)$
$(0,0,25,0,0,3)$ 2.55 2.30 $(0,0,0,0,0,0)$ 0.21 0.17
$(0,0,10,0.5,0.5,1)$
$(0,0,15,1,1,1)$
$(10,10,0,1.5,1.5,2.5)$
Note: $(R_{x},R_{y},R_{z},t_{x},t_{y},t_{z})$ denotes the corresponding rotation and translation components, with units of degrees and meters, respectively.
5.6 Generalizability and Prospective Practical Applications of L2M-Reg
The generalizability of L2M-Reg can be discussed from two perspectives. The first perspective concerns the source of the model data, specifically whether LoD2 models from different agencies can be used. Although L2M-Reg is plane-based and relies on plane information derived from LoD2 models, it is not restricted to models produced by specific agencies or regions. Only the vertex information of each wall surface is required to derive the reference plane parameters. As long as the LoD2 model follows established standards, such as CityGML, and provides access to this information, subsequent processing is independent of the model’s origin.
A fundamental assumption underlying the use of LoD2 models is that building walls are constructed from footprint points that geometrically correspond to the wall plinth. Under this assumption, accurate extraction of the wall plinth area from LiDAR point clouds is essential. This assumption is consistent with common modeling practices across many countries and regions and is therefore considered well-founded. To illustrate this point, the Ingolstadt Store dataset was included as a testbed outside Munich. The LoD2 models and point clouds for this dataset were collected and generated by different agencies than those used for the four Munich datasets, as shown in Table 1. Although the data originate from a different city, all datasets satisfy the footprint point assumption. From a data source perspective, this demonstrates that L2M-Reg is not tied to a specific agency or region and exhibits good generalizability.
The second perspective concerns architectural style, namely whether the method remains applicable to buildings with substantially different designs. This question is closely related to the implementation of the method. L2M-Reg is plane-based, and the algorithm processes facades independently, wall by wall. As a result, its applicability depends on the geometric characteristics of each individual facade. In essence, the key condition is whether a wall surface remains predominantly planar. If this condition is satisfied, this plane-based method remains applicable. If not, the underlying assumption of planarity is violated, and L2M-Reg becomes inapplicable. This limitation is also inherent to many plane-based solutions and represents a clear direction for future improvement.
Regarding the prospective practical applications enabled by L2M-Reg, three key aspects can be summarized:
  1. (1)
    It supports refined, low-cost building-level model updating. It is effective at both the building and the component level. As a result, L2M-Reg substantially reduces data acquisition requirements. High-precision LiDAR-to-Model registration can be achieved by collecting point clouds only in spatially critical areas.
  2. (2)
    It significantly broadens the applicability of LoD2 models in high-precision scenarios. Traditionally, LoD2 models have been used for their large-scale coverage in city-level analyses. Now, L2M-Reg enables accurate LiDAR-to-Model registration. This allows LoD2 models to be reliably used in small-scale, accuracy-sensitive tasks, such as change detection and navigation.
  3. (3)
    It advances the understanding of inherent uncertainty in geospatial data, an aspect often overlooked and considered negligible in LoD2 models. L2M-Reg explicitly connects in-model uncertainty to downstream applications, opening new avenues for uncertainty-aware research and practical applications of geospatial data.